Project

General

Profile

Actions

Ops/Comm Scheduling meeting : 4/4/2017

Location : CR21 in R1 at 14:00 pm

Phone access : New phonebridge number
UK: 08082380274
US: 631.621.5253
Access code: 8161978083
Leader code: 2735

Agenda

AOB:

Minutes

Actions

  • Ajit - No further development of Hanlet SMS system. Need to propose a basic messaging system still to alleviate need for 24Hr monitoring in person. - on hold
    • N.B. that we already have an alarm display in the ISIS MCR. CWhyte to check with ABross/JBoehm that automated texting would be acceptable for removing requirement for continuous oversight while CC magnets powered.
    • We have the SMS to email gateway, including monitoring for e-mail responses, but a lot of work remains in driving this from the Alarm Handler.
  • Macwaters & Boehm - Arrange to install new Sumitomo compressor adsorbers? - stands
  • Macwaters - Re-make dodgy compressor2 remote cable connector - stands
  • Macwaters - Prepare new F70 compressors for use. Remove old overheated CSW71 to R9. Speak to Sumitomo on April 5th regards how to repair. - in progress
    • New F70s on site - RHale to fit mains leads, then move to Hall?
  • Macwaters - Replace old vac sensors and new valve on cryo 1&2 - in progress
  • Ajit - Decide on new way to implement what was MOM/DC state machine override function - in progress
    • Done for DS, needs re-coding for the rest; also need documentation and training for operators.
  • Macwaters - Report on status of FC psu repair (stands)
    • Has arrived at Sorenson and been examined, but no feedback to us as yet
  • Hall Manager - water system change-over procedure (stands)
  • EOverton - Status of miceraid5 disk cleanup (stands)
    • Was done for 2016/05, needs checking for 2017/01
  • SBoyd - ensure MICE-wide discussion of MOM C&M panel superseded
  • MUchida - investigate moving Tracker waveguides around to cure "hole". Is Ed happy for this to progress? Yes. (stands)
  • MCourthold - show JBoehm how to use the Picologger for quench data. Craig can help. (stands)
  • JBoehm & Macwaters - get the West Wall compressors correctly labelled, as various have been re-purposed (stands)
  • CWhyte to check with ABross/JBoehm that automated texting would be acceptable for removing requirement for continuous oversight while CC magnets powered.
  • CWhyte to contact DL to clarify their Hall Task List items
  • PFranchini to go through servers and turn off IPMI in BIOS

DC update (PHodgson)

  • MTucker has swapped turbo pumps between SSU and SSD to reduce wear and tear
  • Preparation for cold-head swap tomorrow: Tracker cryostats have been heated and pumped/purged
  • Neutron monitor: no remote readout, we have a webcam on it instead
  • There is a small leak in the roof water system - about 10%/day
  • Leak checking of the LH2 system is presently in progress; some of this is overseen by Roy Preece and is acting as a de facto review of the documentation

LH2 update (CWhyte)

  • A number of issues have been found, relating to over-pressure bursting windows and putting an asymmetric load on to the FC
    • Putting in a new vent line, on to the spare turret
    • Steve Harrison is double-checking the rate at which LH2 would actually boil off
  • Need another HAZOP - plan is to push ahead as far as possible short of H2 itself
  • Can run with Neon as well, to shake down the system
    • This still has the pressure risk and needs Steve Harrison input - this feedback sets the timetable for closing up the PRY
    • A Neon run would only last about a day

Ops/Commissioning/AOB

Aspirational Pathway: after HAZOP will only reach LH2 late in next cycle. Start earlier with field-off empty-absorber running in evenings, to allow LH2 work to continue during the day.
Once the PRY is closed (see above) can consider testing Diffuser at field and/or investigating the FC quench.

  • MUchida will repeat the leak-stopping on the upstream Tracker butterfly over next couple of days
  • Concern raised about mystery network traffic and unknown devices
    • at least some of this is IPMI boards making DHCP requests
    • PFranchini to go through and disable in BIOS where possible
  • EOverton has got the NI drivers running on replacement target1ctl; need to get RATS running for when Joe comes down prior to BPS test on 20th April.
  • HNebrensky needs to test File Compactor: DATE should allow a Pulser run after Thursday (Tracker fix)
  • AKurup to restart miceecserv after ALH/SM updates - Thursday
  • Surveys:
    • March survey was after PRY opened and FC withdrawn, and SS's moved
    • TOF1 was removed for Diffuser access
    • will need TOF1 and FC re-survey

Updated by Nebrensky, Henry about 6 years ago ยท 6 revisions